OpenStudy (anonymous):

Jason's customed hats experienced fixed cost of $300 and a variable cost of $2 a hat. Write an equation that can be used to determine the total expenses encountered by Jason's custom hats. Let x be the number of hats and c(x) be the total cost of producing x hats. Then c(x)=____? The cost of producing 8 hats is $____?

OpenStudy (anonymous):

c(x) = 300 + 2x Therefore, for 8 hats, the total cost c(x) = 300 + 2*8= 300+16 = 316